# Reducing Optimism Bias in Incomplete Cooperative Games

### v1

@@ D.1. Factory and Supermodular
Finally, we present the study of percentage of selected coalitions for the supermodular(5) class of games. Our results are presented in Figure 7 and 8 All algorithms show preference towards larger coalitions initially. At later steps, PPO deviates from the other strategies and selects coalitions pseudo-uniformly, while the rest favors coalitions of size three. We speculate this is because the utopian gap is very small at this stage, making the actions seem near equal in terms of the reward, i.e. the utopian gap.
@@ Appendix E Oracle algorithms
¯$\overline{}\mathcal{K}\leftarrow 2^{N}\setminus\mathcal{K}_{0}$$\left\{S_{i}\right\}_{i=1}^{t}\leftarrow\operatornamewithlimits{argmin}_{\mathcal{S}\subseteq\overline{}\mathcal{K}:\left\lvert\mathcal{S}\right\rvert=t}\mathcal{G}_{(N,v)}(\mathcal{S})$ return $\{S_{i}\}_{i=1}^{t}$
@@
$\{S_{i}\}_{i=1}^{t-1}\leftarrow$ Oracle Greedy($v,t-1$) ¯$\overline{}\mathcal{K}\leftarrow 2^{N}\setminus(\mathcal{K}_{0}\cup\left\{S_{i}\right\}_{i=1}^{t-1}$) $S_{t}\leftarrow\operatornamewithlimits{argmin}_{S\in\overline{}\mathcal{K}}\mathcal{G}_{(N,v)}(\{S_{i}\}_{i=1}^{t-1}\cup\left\{S\right\},v)$ return $\{S_{i}\}_{i=1}^{t}$
@@ Appendix F Value Function Normalization
Appendix G Example of non-supermodular gap
This section deals with the question of supermodularity of the gap function. We show that as long as $|N|\leq 4$, the gap is supermodular, however, when $|N|\geq 5$, the supermodularity does not have to hold for different subsets of superadditive games as can be seen for the game from Example 3.2.2. We also give examples of totally monotonic (thus also supermodular), symmetric, and graph games. Before we prove Proposition 3, we state several useful lemmata.
@@ Appendix G Example of non-supermodular gap
[…]
@@
[…]
@@
This inequality follows from Lemma G.
@@
    Since $\lvert N\rvert=4$ and $Z,T$ are unknown in $\hat{\mathcal{K}}\cup S$, they must be of form $Z=\{i,j\}$ and $T=\{i,j,k\}$. By Lemma G, $\underline{v}_{\hat{\mathcal{K}}\cup Z}(T)=\underline{v}_{\mathcal{K}}(T)$ and $\underline{v}_{\hat{\mathcal{K}}\cup S\cup Z}(T)=\underline{v}_{\hat{\mathcal{K}}\cup S}(T)$, thus (25) reduces to
@@
    For a contradiction, suppose the converse holds. As, by Lemma G, $\overline{v}_{\hat{\mathcal{K}}\cup S\cup Z}(T)-\overline{v}_{\hat{\mathcal{K}}\cup Z}(T)\geq 0$, this means $\overline{v}_{\hat{\mathcal{K}}\cup Z}(T)>\overline{v}_{\mathcal{K}}(T)$, which leads, by Lemma G, to a contradiction.
@@
    $T\subsetneq Z$: Similarly to the first case, we have $T=\{i,j\}$, $Z=\{i,j,k\}$ and by Lemma G, (25) reduces to
@@
    For a contradiction, suppose the converse holds. It means $\underline{v}_{\mathcal{K}}(T)>\underline{v}_{\hat{\mathcal{K}}\cup Z}(T)$, which by Lemma G leads to a contradiction.
@@
    which holds by Lemma G. In this case, $Z=\{i,j\}$, $T=\{k,\ell\}$ and (25) reduces to
@@
    as $\overline{v}_{\hat{\mathcal{K}}\cup S\cup Z}(T)=\overline{v}_{\hat{\mathcal{K}}\cup S}(T)$ and $\overline{v}_{\hat{\mathcal{K}}\cup Z}(T)=\overline{v}_{\hat{\mathcal{K}}}(T)$. For a contradiction, if $<$ holds, similarly to previous cases, $\underline{v}_{\hat{\mathcal{K}}}(T)>\underline{v}_{\hat{\mathcal{K}}\cup S}(T)$, a contradiction with Lemma G.
@@
    $Z\cap T\neq\emptyset$ and $Z\not\subseteq T$ and $T\not\subseteq Z$: By Lemma G, condition (25) reduces to $0\geq 0$.
G.1. Examples of non-supermodularity
@@
(Game from Example 3.2.2)
For $N=\{1,2,3,4,5\}$ and $o=3$, the condition is violated since,
$\mathcal{G}_{\Delta}=-0.2\not\geq 0.$
Thus, the supermodularity of the gap is violated.
[…]
@@
Any totally monotonic game can be then expressed as $\sum_{S\subseteq N}\alpha_{S}u_{S}$ where $\alpha_{S}\geq 0$ for every $S$.
@@
(Symmetric game) Symmetric cooperative game $(N,v)$ satisfies for every $S$, $T\subseteq N$, $|S|=|T|$ that $v(S)=v(T)$. Symmetric game of $n$ players can be thus defined by $s_{1},\dots,s_{n}\in\mathbb{R}$ corresponding to values of coalitions of different sizes. It holds that if symmetric game on 5 players satisfy $s_{2}=0$, $s_{3}=1$ and $s_{5}=2$, it follows $\mathcal{G}_{\Delta}=-0.1\not\geq 0$.
@@
For a complete graph $(N,E)$, where $N=\{1,2,3,4,5\}$, $w_{1,2}=1$, $w_{2,3}=4$, and $w_{3,4}=5$, it holds $\mathcal{G}_{\Delta}=-0.1\not\geq 0$.
Appendix H Explicit form of the gap
$\mathcal{G}_{(N,v)}(\mathcal{K})=\sum_{i\in N}\sum_{S\subseteq N\setminus i}\beta_{S}\left(\overline{v}_{\hat{\mathcal{K}}}(S\cup i)-\underline{v}_{\hat{\mathcal{K}}}(S)\right)-v(N),$ (26)
where $\beta_{S}=\frac{|S|!(|N|-|S|-1)!}{|N|!}$. We rewrite this double sum as a sum of $S\subseteq N$. First, $\overline{v}_{\hat{\mathcal{K}}}(S)$ appears in (26) for every $i\in S$ with coefficient $\beta_{S\setminus i}$. Second, $\underline{v}_{\hat{\mathcal{K}}}(S)$ appears in (26) for every $i\in N\setminus S$ with coefficient $\beta_{S}$. We get for any $i\in S$,
$\mathcal{G}_{(N,v)}(\mathcal{K})=\sum_{S\subseteq N}\left(\gamma_{S}\overline{v}_{\hat{\mathcal{K}}}(S)-\delta_{S}\underline{v}_{\hat{\mathcal{K}}}(S)\right)-v(N),$
where
  • $\gamma_{S}=|S|\frac{(|S|-1)!(|N|-|S|)!}{|N|!}=\frac{|S|!(|N|-|S|)!}{|N|!}$,
  • $\delta_{S}=(|N|-|S|)\frac{|S|!(|N|-|S|-1)!}{|N|!}=\frac{|S|!(|N|-|S|)!}{|N|!}$.
This concludes the proof of Proposition 3, as
$\mathcal{G}_{(N,v)}(\mathcal{K})=\sum_{S\subseteq N}\frac{|S|!(|N|-|S|)!}{|N|!}\left(\overline{v}_{\hat{\mathcal{K}}}(S)-\underline{v}_{\hat{\mathcal{K}}}(S)\right)-v(N).$

### v3

@@ D.1. Factory and Supermodular
Figure 6. The utopian gap as a function of number of revealed coalitions (i.e. steps of the Principals problem) for different sizes. Note that the PPO algorithm was trained on the same number of time steps as for the 4 player problem. Even though the action space size has increased dramatically (it grows exponentially with $n$), it still beats the trivial random approach. Better performance of the PPO approach can still be reached by increasing the learning period. Note also, that due to time constraints, when gathering the data for the plots, we used 100 samples for sizes 8 and 9, and 10 samples for size 10, while also dropping the PPO, as the training process was too time demanding.
@@
Finally, we present the study of percentage of selected coalitions for the supermodular(5) class of games. Our results are presented in Figure 8 and 9 All algorithms show preference towards larger coalitions initially. At later steps, PPO deviates from the other strategies and selects coalitions pseudo-uniformly, while the rest favors coalitions of size three. We speculate this is because the utopian gap is very small at this stage, making the actions seem near equal in terms of the reward, i.e. the utopian gap.
@@ Appendix E Oracle algorithms
1 ¯$\overline{}\mathcal{K}\leftarrow 2^{N}\setminus\mathcal{K}_{0}$
2 $\left\{S_{i}\right\}_{i=1}^{t}\leftarrow\operatornamewithlimits{argmin}_{\mathcal{S}\subseteq\overline{}\mathcal{K}:\left\lvert\mathcal{S}\right\rvert=t}\mathcal{G}_{(N,v)}(\mathcal{S})$
3 return $\{S_{i}\}_{i=1}^{t}$
@@
1 $\{S_{i}\}_{i=1}^{t-1}\leftarrow$ Oracle Greedy($v,t-1$)
2 ¯$\overline{}\mathcal{K}\leftarrow 2^{N}\setminus(\mathcal{K}_{0}\cup\left\{S_{i}\right\}_{i=1}^{t-1}$)
3 $S_{t}\leftarrow\operatornamewithlimits{argmin}_{S\in\overline{}\mathcal{K}}\mathcal{G}_{(N,v)}(\{S_{i}\}_{i=1}^{t-1}\cup\left\{S\right\},v)$
4 return $\{S_{i}\}_{i=1}^{t}$
@@ Appendix F Value Function Normalization
Appendix G Explicit form of the gap
$\mathcal{G}_{(N,v)}(\mathcal{K})=\sum_{i\in N}\sum_{S\subseteq N\setminus i}\beta_{|S|}\left(\overline{v}_{\hat{\mathcal{K}}}(S\cup i)-\underline{v}_{\hat{\mathcal{K}}}(S)\right)-v(N),$ (24)
where $\beta_{|S|}=\frac{|S|!(|N|-|S|-1)!}{|N|!}$. We rewrite this double sum as a sum of $S\subseteq N$. First, $\overline{v}_{\hat{\mathcal{K}}}(S)$ appears in (24) for every $i\in S$ with coefficient $\beta_{|S|-1}$. Second, $\underline{v}_{\hat{\mathcal{K}}}(S)$ appears in (24) for every $i\in N\setminus S$ with coefficient $\beta_{|S|}$. We get for any $i\in S$,
$\mathcal{G}_{(N,v)}(\mathcal{K})=\sum_{S\subseteq N}\left(\gamma_{|S|}\overline{v}_{\hat{\mathcal{K}}}(S)-\delta_{|S|}\underline{v}_{\hat{\mathcal{K}}}(S)\right)-v(N),$
where
  • $\gamma_{|S|}=|S|\frac{(|S|-1)!(|N|-|S|)!}{|N|!}=\frac{|S|!(|N|-|S|)!}{|N|!}$,
  • $\delta_{|S|}=(|N|-|S|)\frac{|S|!(|N|-|S|-1)!}{|N|!}=\frac{|S|!(|N|-|S|)!}{|N|!}$.
Notice there are two edge-cases, where the second equality does not hold:
  1. (1)
    If $S=\emptyset$, then $\gamma_{0}=0$, which is lost when $S$ joins the factorial expression, yielding $0!=1$. However, $\gamma_{0}$ is always multiplied by $\left(\overline{v}_{\hat{}\mathcal{K}}(S)-\underline{v}_{\hat{}\mathcal{K}}(S)\right)=\left(0-0\right)=0$, so the value of this coefficient does not change the resulting value.
  2. (2)
    If $S=N$, then $\gamma_{n}=0$, however, $\frac{|S|!(|N|-|S|)!}{|N|!}=1$. In fact, $\overline{v}_{\hat{}\mathcal{K}}(S)$ appears exactly once in the sum, but $\underline{v}_{\hat{}\mathcal{K}}(S)$ does not appear at all. Our original calculations, however, didn’t take into account the final $-v(N)$ term present at the end of the definition of the gap. With the subtraction of $v(N)=\underline{v}_{\hat{}\mathcal{K}}(S)$, it is easy to see that the value of $\gamma_{n}=1$ is correct.
This concludes the proof of Proposition 3, as
$\mathcal{G}_{(N,v)}(\mathcal{K})=\sum_{S\subseteq N}\frac{|S|!(|N|-|S|)!}{|N|!}\left(\overline{v}_{\hat{\mathcal{K}}}(S)-\underline{v}_{\hat{\mathcal{K}}}(S)\right).$
Appendix H Supermodularity of gap
Supermodularity of the utopian gap is a desirable property, as it is known that greedy algorithms perform well on supermodular function (George et al., 1978, Proposition 3.4). We show that as long as $|N|\leq 4$, the gap is supermodular, however, when $|N|\geq 5$, the supermodularity does not have to hold for different subsets of superadditive games. For $|N|=5$, we give examples from different important subsets of superadditive games that do not satisfy supermodularity of the gap. Among these are totally monotonic (thus also supermodular), symmetric, and graph games. For $|N|\geq 6$, we derive a criterion, which, when satisfied, implies non-supermodularity of the gap. As a corollary, we show all three of the above mentioned classes contain games with non-supermodular gap.
Throughout this section, we implicitly assume that the value function $v:2^{N}\to\mathbb{R}$ is normalized as described in Appendix F. This is without the loss of generality, as this preserves the value of the gap, see Lemma F. As a consequence, the lower function $\underline{v}(S)$ is zero for $\mathcal{K}=\mathcal{K}_{0}$ for any $S\in 2^{N}\setminus\mathcal{K}_{0}$.
H.1. Supermodularity for $|N|\leq 4$
Before we prove Proposition 3.2.2, we state several useful lemmata.
@@ Appendix H Supermodularity of gap
$\underline{v}_{\hat{\mathcal{K}}}(T)\leq\underline{v}_{\hat{\mathcal{K}}\cup S}(T)\hskip 8.61108pt\text{and}\hskip 8.61108pt\overline{v}_{\hat{\mathcal{K}}}(T)\geq\overline{v}_{\hat{\mathcal{K}}\cup S}(T).$
@@
$\overline{v}_{\hat{\mathcal{K}}}(T)=\overline{v}_{\hat{\mathcal{K}}\cup S}(T),\hskip 12.91663pt\text{and}\hskip 12.91663pt\underline{v}_{\hat{\mathcal{K}}}(T)=\underline{v}_{\hat{\mathcal{K}}\cup S}(T).$
@@
This inequality follows from Lemma H.1.
@@
    Since $\lvert N\rvert=4$ and $Z,T$ are unknown in $\hat{\mathcal{K}}\cup S$, they must be of form $Z=\{i,j\}$ and $T=\{i,j,k\}$. By Lemma H.1, $\underline{v}_{\hat{\mathcal{K}}\cup Z}(T)=\underline{v}_{\mathcal{K}}(T)$ and $\underline{v}_{\hat{\mathcal{K}}\cup S\cup Z}(T)=\underline{v}_{\hat{\mathcal{K}}\cup S}(T)$, thus (25) reduces to
@@
    For a contradiction, suppose the converse holds. As, by Lemma H.1, $\overline{v}_{\hat{\mathcal{K}}\cup S\cup Z}(T)-\overline{v}_{\hat{\mathcal{K}}\cup Z}(T)\geq 0$, this means $\overline{v}_{\hat{\mathcal{K}}\cup Z}(T)>\overline{v}_{\mathcal{K}}(T)$, which leads, by Lemma H.1, to a contradiction.
@@
    $T\subsetneq Z$: Similarly to the first case, we have $T=\{i,j\}$, $Z=\{i,j,k\}$ and by Lemma H.1, (25) reduces to
@@
    For a contradiction, suppose the converse holds. It means $\underline{v}_{\mathcal{K}}(T)>\underline{v}_{\hat{\mathcal{K}}\cup Z}(T)$, which by Lemma H.1 leads to a contradiction.
@@
    which holds by Lemma H.1. In this case, $Z=\{i,j\}$, $T=\{k,\ell\}$ and (25) reduces to
@@
    as $\overline{v}_{\hat{\mathcal{K}}\cup S\cup Z}(T)=\overline{v}_{\hat{\mathcal{K}}\cup S}(T)$ and $\overline{v}_{\hat{\mathcal{K}}\cup Z}(T)=\overline{v}_{\hat{\mathcal{K}}}(T)$. For a contradiction, if $<$ holds, similarly to previous cases, $\underline{v}_{\hat{\mathcal{K}}}(T)>\underline{v}_{\hat{\mathcal{K}}\cup S}(T)$, a contradiction with Lemma H.1.
@@
    $Z\cap T\neq\emptyset$ and $Z\not\subseteq T$ and $T\not\subseteq Z$: By Lemma H.1, condition (25) reduces to $0\geq 0$.
H.2. Non-supermodularity for $|N|=5$
@@
Any totally monotonic game can be then expressed as $\sum_{S\subseteq N}\alpha_{S}u_{S}$ where $\alpha_{S}\geq 0$ for every $S$. They form a subset of convex games, which under the name belief measures forms the foundation on the evidence theory Grabisch (2016); Shafer (1976))
@@
(Symmetric game) Symmetric cooperative game $(N,v)$ satisfies for every $S$, $T\subseteq N$, $|S|=|T|$ that $v(S)=v(T)$. Symmetric game of $n$ players can be thus defined by $s_{1},\dots,s_{n}\in\mathbb{R}$ corresponding to values of coalitions of different sizes. These games are often studied for their robust properties with respect to solution concepts Peleg and Sudhölter (2007)).
It holds that if symmetric game on 5 players satisfy $s_{2}=s_{3}=1$ and $s_{5}=2$, it follows $\mathcal{G}_{\Delta}=-0.1\not\geq 0$.
@@
For a complete graph $(N,E)$, where $N=\{1,2,3,4,5\}$, $w_{1,2}=1=w_{3,4}=1$, it holds $\mathcal{G}_{\Delta}=-0.1\not\geq 0$.
H.3. Criterion for $|N|\geq 6$
Proposition 3.2.2.
For $|N|\geq 6$, the gap of an incomplete $\mathbb{S}^{n}$-extendable game $(N,w)$ is not supermodular if there exist $i,j,k,l\in N$ such that $v(ij)\leq v(jk)\leq v(kl)$, where $v(S)=w(S)-\sum_{i\in S}w(\{i\})$, and
$v(kl)<\left[{\binom{n}{2}}{\binom{n}{\lfloor n/2\rfloor}}^{-1}\left(2^{n-3}-n+2\right)-1\right]v(ij).$ (26)
Proof.
By Lemma F in Appendix F, the gap of $(N,w)$ is supermodular if and only if the gap of $(N,v)$ is supermodular. We focus on the gap $(N,v)$ as this makes our analysis simple and at the same time, the condition easy to check.
The supermodularity condition (3) for the gap is
$\mathcal{G}_{(N,v)}(\hat{\mathcal{K}}\cup S\cup Z)-\mathcal{G}_{(N,v)}(\hat{\mathcal{K}}\cup S)\geq\mathcal{G}_{(N,v)}(\hat{\mathcal{K}}\cup Z)-\mathcal{G}_{(N,v)}(\hat{\mathcal{K}}),$ (27)
where $\hat{\mathcal{K}}\subseteq 2^{N}\setminus\mathcal{K}_{0}$.
We have shown in Appendix G that the utopian gap is
$\mathcal{G}_{(N,v)}(\mathcal{K})=\sum_{S\subseteq N}\gamma_{S}\left(\overline{v}_{\hat{\mathcal{K}}}(S)-\underline{v}_{\hat{\mathcal{K}}}(S)\right)-v(N),$ (28)
where $\gamma_{S}=\frac{|S|!(|N|-|S|)!}{|N|!}$. This equation can be viewed as a weighted $l_{1}$-norm of the difference between the upper and lower games. [^6] Assuming they are viewed as vectors in $\mathbb{R}^{2^{n}}$.
Using this equality, we can rewrite Eq. (27) as
$\sum_{T\subseteq N}\gamma_{T}\left(\overline{v}_{\hat{\mathcal{K}}\cup S\cup Z}(T)-\underline{v}_{\hat{\mathcal{K}}\cup S\cup Z}(T)-\overline{v}_{\hat{\mathcal{K}}\cup S}(T)+\underline{v}_{\hat{\mathcal{K}}\cup S}(T)\right)\geq\\ \sum_{T\subseteq N}\gamma_{T}\left(\overline{v}_{\hat{\mathcal{K}}\cup Z}(T)-\underline{v}_{\hat{\mathcal{K}}\cup Z}(T)-\overline{v}_{\hat{\mathcal{K}}}(T)+\underline{v}_{\hat{\mathcal{K}}}(T)\right).$ (29)
We obtain the result by setting $\mathcal{K}=\mathcal{K}_{0}\cup\{\{j,k\}\}$, $S=\{i,j\}$, and $Z=\{k,l\}$. Denote $v(ij)=v(jk)-\varepsilon$ and $v(kl)=v(jk)+\delta$ for some $\varepsilon,\delta\geq 0$. For these values, we investigate each of the weighted terms separately. Let
$L_{T}$ $=\overline{v}_{\hat{\mathcal{K}}\cup S\cup Z}(T)-\underline{v}_{\hat{\mathcal{K}}\cup S\cup Z}(T)-\overline{v}_{\hat{\mathcal{K}}\cup S}(T)+\underline{v}_{\hat{\mathcal{K}}\cup S}(T),$ (30)
$R_{T}$ $=\overline{v}_{\hat{\mathcal{K}}\cup Z}(T)-\underline{v}_{\hat{\mathcal{K}}\cup Z}(T)-\overline{v}_{\hat{\mathcal{K}}}(T)+\underline{v}_{\hat{\mathcal{K}}}(T).$ (31)
In the reminder of the proof, we distinguish different cases based on the relation of $T$ and $\{i,j,k,l\}$ for all $T\in 2^{N}\setminus\mathcal{K}_{0}$.
  1. (1)
    $\{i,j,k,l\}\subseteq T\subsetneq N$:
    Using Lemma H.1, we get
    $L_{T}$ $=\underline{v}_{\mathcal{K}\cup S}(T)-\underline{v}_{\mathcal{K}\cup S\cup Z}(T),$
    $R_{T}$ $=\underline{v}_{\mathcal{K}}(T)-\underline{v}_{\mathcal{K}\cup Z}(T),$
    which can be expressed as
    $L_{T}$ $=v(jk)-v(jk)+\varepsilon-v(jk)-\delta,$
    $R_{T}$ $=v(jk)-v(jk)-\delta.$
    It holds $L_{T}=R_{T}+\varepsilon-v(jk)=R_{T}-v(ij)$.
  2. (2)
    $\{i,j,k,l\}\subseteq N\setminus T$:
    In this case, the lower bound on $v(T)$ is unaffected by the knowledge of $S,Z$, so
    $L_{T}$ $=\overline{v}_{\mathcal{K}\cup S\cup Z}(T)-\overline{v}_{\mathcal{K}\cup S}(T),$
    $R_{T}$ $=\overline{v}_{\mathcal{K}\cup Z}(T)-\overline{v}_{\mathcal{K}}(T).$
    Since the only superset of $T$ in $\mathcal{K}\cup S\cup Z$ is $N$, the upper bound (9) reduces to
    $\overline{v}_{\mathcal{L}}(T)=v(N)-\underline{v}_{\mathcal{L}}(N\setminus T)=1-\underline{v}_{\mathcal{L}}(N\setminus T),$
    for every $\mathcal{L}\in\{\mathcal{K},\mathcal{K}\cup S,\mathcal{K}\cup Z,\mathcal{K}\cup S\cup Z\}$. Thus
    $L_{T}$ $=\underline{v}_{\mathcal{K}\cup S}(N\setminus T)-\underline{v}_{\mathcal{K}\cup S\cup Z}(N\setminus T),$
    $R_{T}$ $=\underline{v}_{\mathcal{K}}(N\setminus T)-\underline{v}_{\mathcal{K}\cup Z}(N\setminus T).$
    Further, $i,j,k,l\in N\setminus T$, which means
    $L_{T}$ $=v(jk)-v(jk)+\varepsilon-v(jk)-\delta,$
    $R_{T}$ $=v(jk)-v(jk)-\delta.$
    As in the previous case, $L_{T}=R_{T}+\varepsilon-v(jk)=R_{T}-v(ij)$.
  3. (3)
    $S=T$:
    In this case, since $\underline{v}_{\mathcal{L}}(T)=\overline{v}_{\mathcal{L}}(T)=v(T)$ if $T\in\mathcal{L}$ for every $\mathcal{L}\in\{\mathcal{K},\mathcal{K}\cup S,\mathcal{K}\cup Z,\mathcal{K}\cup S\cup Z\}$ we get
    $L_{T}$ $=0,$
    $R_{T}$ $=\overline{v}_{\mathcal{K}\cup Z}(T)-\overline{v}_{\mathcal{K}}(T)=\underline{v}_{\mathcal{K}}(N\setminus T)-\underline{v}_{\mathcal{K}\cup Z}(N\setminus T).$
    This means $R_{T}=0-v(jk)-\delta$, therefore it holds $L_{T}=R_{T}+v(jk)+\delta=R_{T}+v(kl)$.
  4. (4)
    $Z=T$:
    In this case, by the fact that $\underline{v}_{\mathcal{L}}(T)=\overline{v}_{\mathcal{L}}(T)=v(T)$ if $T\in\mathcal{L}$, it holds
    $L_{T}$ $=\underline{v}_{\mathcal{K}\cup S}(T)-\overline{v}_{\mathcal{K}\cup S}(T),$
    $R_{T}$ $=\underline{v}_{\mathcal{K}}(T)-\overline{v}_{\mathcal{K}}(T).$
    Since $S\not\subseteq T$, it holds $\underline{v}_{\mathcal{K}\cup S}(T)=\underline{v}_{\mathcal{K}}(T)$, so we can rewrite
    $L_{T}$ $=\underline{v}_{\mathcal{K}\cup S}(N\setminus T)-v(N)=v(jk)-\varepsilon-v(N),$
    $R_{T}$ $=\underline{v}_{\mathcal{K}}(N\setminus T)-v(N)=-v(N).$
    Thus it holds $L_{T}=R_{T}+v(jk)-\varepsilon=R_{T}+v(ij)$.
  5. (5)
    $\emptyset\neq Z\cap T,Z\not\subseteq T$, and $T\not\subseteq Z$ or $\emptyset\neq S\cap T,S\not\subseteq T$, and $T\not\subseteq S$:
    By Lemma H.1, and Lemma H.1, one immediately arrives at $L_{T}=R_{T}=0$.
A necessary condition for supermodularity of the gap is
$\sum_{T\subseteq N}\gamma_{|T|}(L_{T}-R_{T})\geq 0.$
This means, that if any upper bound $L$,
$\sum_{T\subseteq N}\gamma_{|T|}(L_{T}-R_{T})\leq L$
satisfies $L<0$, the gap is not supermodular. To construct the upper bound, we focus on terms of the sum corresponding to those from the first and the second case above, i.e. $T\subseteq N$ s.t. $\{i,j,k,l\}\subseteq T\subsetneq N$ or $T\subseteq N\setminus\{i,j,k,l\}$. In these cases, $L_{T}-R_{T}\leq 0$. Since $\gamma_{|T|}\leq\gamma_{\lfloor n/2\rfloor}$ for every $T\subseteq N$, we bound the corresponding terms by
$\gamma_{|T|}\left(L_{T}-R_{T}\right)\leq\gamma_{\lfloor n/2\rfloor}\left(L_{T}-R_{T}\right).$
The upper bound thus sums to
$L=-\gamma_{\lfloor n/2\rfloor}(c_{1}(n)+c_{2}(n))v(ij)+\gamma_{2}(v(kl)+v(ij)),$
where $c_{1}(n)=2^{n-4}-1$ and $c_{2}(n)=2^{n-4}-n+3$ is the number of terms satisfying conditions 1 and 2, respectively. Thus, $L<0$ is equal to
$v(kl)<\left(\frac{\gamma_{\lfloor n/2\rfloor}}{\gamma_{2}}\left(c_{1}(n)+c_{2}(n)\right)-1\right)v(ij),$
or since $c_{1}(n)+c_{2}(n)=2^{n-3}-n+2$, it is equal to
$v(kl)<\left[{\binom{n}{2}}{\binom{n}{\lfloor n/2\rfloor}}^{-1}\left(2^{n-3}-n+2\right)-1\right]v(ij).$ (32)
The coefficient for $v(ij)$ is equal $2,\frac{28}{5},\frac{47}{5}$$\dots$ for $n=6,7,8,\dots$ and the sequence can be shown to be increasing. For $n=5$, the coefficient is equal to 00, which makes the criterion useless as it translates to $00\leq v(kl)<$ 0. ∎
A corollary of the criterion is that, in any game where $|N|\geq 6$ and there exist $i,j,k,l\in N$ such that $v(ij)\leq v(jk)\leq v(kl)$ and $v(kl)<2v(ij)$ does not satisfy supermodularity of the gap. The criterion is satisfied by any game satisfying $v(ij)=v(jk)=v(kl)\neq 0$ for some $i,j,k,l$. Examples of totally monotonic, symmetric supperadditive, and graph games are immediate. Note that almost every symmetric superadditive ($v(ij)\neq 0$) as well as almost every additive game (different from zero game) has a non-supermodular gap.
H.4. Simpler criterion for $|N|\geq 7$
Using the same technique as in H.3, one might derive weaker, however, simpler criterion. E.g. since $-\gamma_{\lfloor n/2\rfloor}(c_{1}(n)+c_{2}(n))\leq-\frac{1}{10}$ for every $n\geq 5$, it holds $L\leq-\frac{1}{10}v(ij)+\gamma_{2}\left(v(kl)+v(ij)\right)$, which is strictly smaller than 00 if $v(kl)<\left(1-10\gamma_{2}\right)v(ij)/10\gamma_{2}$, or
$v(kl)<\frac{n^{2}-n-20}{20}v(ij).$ (33)
However, this criterion is useful only when $n\geq 7$, as for $n=5,6$, it holds $\alpha_{n}=\frac{n^{2}-n-20}{20}<1$, which implies
$0\leq v(ij)\leq v(kl)<\alpha_{n}v(ij),$
thus $v(ij)=v(kl)=0$ and the criterion simplifies to $00<\alpha_{n}$ 0, which cannot be satisfied. However, for $n\geq 7$, it holds
$\frac{5}{4}v(ij)\leq\alpha_{n}v(ij),$
thus any game satisfying
$0<v(ij)\leq v(kl)<\frac{5}{4}v(ij)$
cannot have a supermodular gap.